# PyTrim: A Practical Tool for Reducing Python Dependency Bloat

### v1

@@ Abstract
Dependency bloat is a persistent challenge in Python projects, which increases maintenance costs and security risks. While numerous tools exist for detecting unused dependencies in Python, removing these dependencies across the source code and configuration files of a project requires manual effort and expertise. To tackle this challenge we introduce pytrim, an end-to-end system to automate this process. pytrim eliminates unused imports and package declarations across a variety of file types, including Python source and configuration files such as requirements.txt and setup.py. pytrim’s modular design makes it agnostic to the source of dependency bloat information, enabling integration with any detection tool. Beyond its contribution when it comes to automation, pytrim also incorporates a novel dynamic analysis component that improves dependency detection recall. Our evaluation of pytrim’s end-to-end effectiveness on a ground-truth dataset of 37 merged pull requests from prior work, shows that pytrim achieves 98.3% accuracy in replicating human-made changes. To show its practical impact, we run pytrim on 971 open-source packages, identifying and trimming bloated dependencies in 39 of them. For each case, we submit a corresponding pull request, 6 of which have already been accepted and merged. pytrim is available as an open-source project, encouraging community contributions and further development.
@@ I Introduction
Python has emerged as one of the most widely adopted programming languages globally [1, 2] partly due to its large ecosystem of third-party packages, with the Python Package Index (PyPI) – currently hosting over 654,000 projects [3]. While these packages accelerate development and foster reuse they also contribute to a notable software engineering challenge: dependency bloat. As projects evolve, obsolete dependencies are often left in the codebase. This creates technical debt, which in turn leads to increased maintenance overhead, expanded attack surface, and slower pipelines [4, 5, 6, 7].
@@
The Python ecosystem provides numerous open-source tools that focus on bloat detection. Some of these tools operate at the file level, identifying unused imports within individual Python files [9, 10]. Others operate at the project level, looking for unused package dependencies by comparing configuration files against the imports used in the source code [11, 12]. While valuable for detection, these tools do not provide a removal solution.
@@
  • Automation of the maintenance lifecycle by creating a new Git branch, committing the changes, and generating a pull request with a detailed report, enabling a “review-and-merge” approach to dependency cleanup.
@@ II Background & Motivation
Once unused dependencies are identified, developers must undertake a manual removal process that involves removing dependency declarations from configuration files and imports from Python files, and finally submitting a pull request to fix the problem. This remediation workflow requires careful attention to ensure that all references to the unused dependency are properly eliminated from the codebase without affecting the package’s functionality.
Challenge 1: Project dependency resolution: To compute the dependency set, existing detection tools rely on static analyses of configuration files. While tools such as fawltydeps, correctly parse standard configuration files such as the ones depicted in Fig. 1, they fail when dependencies are programmatically constructed. For example, in the optimizely-sdk package (Fig. 2), the setup.py file dynamically builds the install_requires list by reading an external file, making static analysis approximation difficult and causing missed dependencies. Given that over 50% of PyPI packages published in April 2024 [15] use setup.py as their configuration file, the practical approximation of its behavior has become essential.
Challenge 2: Multitude of configuration files and types: To remove dependency declarations developers have to search for and edit all related configuration files (e.g. all four aforementioned configuration files in the case of softlayer-python). Notably, open-source projects have previously encountered issues [16] with partial updates of configuration files, leading to broken builds.
@@ III-A Overview
pytrim receives as input the path to a Python project and then proceeds to (a) dynamically infer the project’s dependencies by installing it in isolation (b) identify the subset of unused dependencies by invoking an unused dependency detector, and (c) locate and remove their occurrences (declarations and imports) across a wide range of file formats. Its design is detector-agnostic, allowing any dependency analysis tool to be integrated.
@@ III-C Implementation and Usage
pytrim is implemented in $\sim$4,200 lines of Python code and is provided as a command-line tool, installable from PyPI via pip install pytrim and accessible via the pytrim command. Apart from end-to-end automated analysis (Challenge 3), it can also work in removal-only mode, bypassing its built-in detectors, accepting the list of unused packages as input alongside the project’s path. Further, pytrim can also generate markdown reports and pull requests.
@@ IV-A Dynamic Dependency Resolution
We compare pytrim’s dependency resolver with the static resolvers employed by the three integrated detection tools, to quantify the missed dependencies it helps uncover. To this end, we evaluate each resolver on a dataset of 1300 popular GitHub projects studied in previous work [17]. From these, we only keep a subset of 971 packages that are successfully installed. Compared to the static resolvers, the dynamic resolver is able to uncover missed dependencies in 48/971 (5%) of cases. The main causes for the static resolvers’ misses are: (a) parsing limitations for edge-cases of supported declarative files, such as pyproject.toml (18 cases), (b) arbitrary setup.py code, such as function invocations (22 cases), and (c) miscellaneous bugs (8 cases). For example, the dependency of optimizely-sdk on pyrsistent, shown in Fig. 2, is detected only through our dynamic resolution technique.
@@ IV-B Remover effectiveness
To validate pytrim’s effectiveness, we design an automated pipeline to measure its accuracy against real-world software maintenance tasks. As a ground truth, we use the dataset from Drosos et al. [17], which provides a curated set of pull requests (PRs) where bloated dependencies were manually removed from open-source Python projects.
@@ IV-C Real-World Deployment
To evaluate pytrim’s practical applicability, we deploy it on the 971 installable open-source Python projects studied in previous work [17]. For each project, we run pytrim using the call-graph-based detector to identify unused dependencies. Then, we manually verify all changes and submit pull requests to the respective repositories.
pytrim removes unused dependencies in 39 projects, resulting in 39 submitted pull requests. At the time of writing, 6 of these PRs have been merged by the project maintainers, 29 are still under review, and 4 were closed due to project-specific policies.
As an example consider the softlayer-python project (Fig. 1). pytrim correctly identified that the prettytable dependency is unused. Then, it generated a PR to remove the dependency from four configuration files (setup.py and three separate requirements.txt files). pytrim also identified the unused package name in the README.rst file and correctly flagged it for manual review. The resulting PR, with the manual edit for the README, was merged by the project maintainers. All merged PRs are summarized in Table II.
@@
TABLE II: Merged pull requests contributed by pytrim.
Project PR ID Files Modified
pytroll/pyresample #669 1
furlongm/patchman #689 1
pycqa/prospector #781 2
jamesoff/simplemonitor #1646 4
softlayer/softlayer-python #2230 5
napalm-automation/napalm #2243 1
@@ V Related Work
Existing Python tools focus primarily on detection and can be broadly categorized into two groups. The first consists of source code linters, such as autoflake [9] and pylint [10], that identify unused imports directly within .py files. While some of these, such as autoflake, can remove the identified import lines, their scope is strictly limited to the source code; they do not modify project-level configuration files. The second group operates at the project level; tools such as deptry [12], and FawltyDeps [11] compare declared dependencies against source code imports to find unused packages, but they do not automate the removal of the corresponding import statements.

### v2

@@ Abstract
Dependency bloat is a persistent challenge in Python projects, which increases maintenance costs and security risks. While numerous tools exist for detecting unused dependencies in Python, removing these dependencies across the source code and configuration files of a project requires manual effort and expertise. To tackle this challenge we introduce pytrim, an end-to-end system to automate this process. pytrim eliminates unused imports and package declarations across a variety of file types, including Python source and configuration files such as requirements.txt and setup.py. pytrim’s modular design makes it agnostic to the source of dependency bloat information, enabling integration with any detection tool. Beyond its contribution when it comes to automation, pytrim also incorporates a novel dynamic analysis component that improves dependency detection recall. Our evaluation of pytrim’s end-to-end effectiveness on a ground-truth dataset of 37 merged pull requests from prior work, shows that pytrim achieves 98.3% accuracy in replicating human-made changes. To show its practical impact, we run pytrim on 971 open-source packages, identifying and trimming bloated dependencies in 39 of them. For each case, we submit a corresponding pull request, 14 of which have already been accepted and merged. pytrim is available as an open-source project, encouraging community contributions and further development.
@@ I Introduction
Python has emerged as one of the most widely adopted programming languages globally [1, 2] partly due to its large ecosystem of third-party packages, with the Python Package Index (PyPI) – currently hosting over 654,000 projects [3]. While these packages accelerate development and foster reuse they also contribute to a notable software engineering challenge: dependency bloat. As projects evolve, obsolete dependencies are often left in the codebase. This creates technical debt, which in turn leads to increased maintenance overhead, expanded attack surface (e.g., via __init__.py, which is executed on import), and slower pipelines [4, 5, 6, 7].
@@
The Python ecosystem provides numerous open-source tools that focus on bloat detection. Some operate at the file level, identifying unused imports within individual Python files [9, 10]. Others operate at the project level, looking for unused package dependencies by comparing configuration files against the imports used in the source code [11, 12]. While valuable for detection, these tools do not provide a removal solution.
@@
  • Automation of the maintenance lifecycle by creating a new Git branch, committing the changes, and generating a pull request (PR) with a detailed report, enabling a “review-and-merge” approach to dependency cleanup.
@@ II Background & Motivation
Once unused dependencies are identified, developers must undertake a manual removal process that involves removing dependency declarations from configuration files and imports from Python files, and finally submitting a PR to fix the problem. This remediation workflow requires careful attention to ensure that all references to the unused dependency are properly eliminated from the codebase without affecting the package’s functionality.
Challenge 1: Project dependency resolution: To compute the dependency set, existing detection tools rely on static analysis of configuration files. While tools such as FawltyDeps, correctly parse standard configuration files such as the ones depicted in Fig. 1, they fail when dependencies are programmatically constructed. For example, in the optimizely-sdk package (Fig. 2), the setup.py file dynamically builds the install_requires list by reading an external file, making static analysis approximation difficult and causing missed dependencies. Given that over 50% of PyPI packages published in April 2024 [15] use setup.py as their configuration file, the practical approximation of its behavior is essential.
Challenge 2: Multitude of configuration files and types: To remove dependency declarations developers have to search for and edit all related configuration files (e.g., all four aforementioned configuration files in the case of softlayer-python). Notably, open-source projects have previously encountered issues [16] with partial updates of configuration files, leading to broken builds.
@@ III-A Overview
pytrim receives as input the path to a Python project and then proceeds to (a) dynamically infer the project’s dependencies by installing it in isolation, (b) identify the subset of unused dependencies by invoking an unused dependency detector, and (c) locate and remove their occurrences (declarations and imports) across a wide range of file formats. Its design is detector-agnostic, allowing any dependency analysis tool to be integrated.
@@ III-C Implementation and Usage
pytrim is implemented in $\sim$4,500 lines of Python code and is provided as a command-line tool, installable from PyPI via pip install pytrim and accessible via the pytrim command. Apart from end-to-end automated analysis (Challenge 3), it can also work in removal-only mode, bypassing its built-in detectors, accepting the list of unused packages as input alongside the project’s path. Further, pytrim can also generate markdown reports and PRs.
@@ IV-A Dynamic Dependency Resolution
We compare pytrim’s dependency resolver with the static resolvers employed by the three integrated detection tools, to quantify the missed dependencies it helps uncover. To this end, we evaluate each resolver on a dataset of 1,300 popular GitHub projects studied in previous work [17]. From these, we only keep a subset of 971 packages that are successfully installed. Compared to the static resolvers, the dynamic resolver is able to uncover missed dependencies in 48/971 (5%) of cases. The main causes for the static resolvers’ misses are: (a) parsing limitations for edge-cases of supported declarative files, such as pyproject.toml (18 cases), (b) arbitrary setup.py code, such as function invocations (22 cases), and (c) miscellaneous bugs (8 cases). For example, the dependency of optimizely-sdk on pyrsistent, shown in Fig. 2, is detected only through our dynamic resolution technique.
@@ IV-B Remover effectiveness
To validate pytrim’s effectiveness, we design an automated pipeline to measure its accuracy against real-world software maintenance tasks. As a ground truth, we use the dataset from Drosos et al. [17], which provides a curated set of PRs where bloated dependencies were manually removed from open-source Python projects.
@@ IV-C Real-World Deployment
To evaluate pytrim’s practical applicability, we deploy it on the 971 installable open-source Python projects studied in previous work [17]. For each project, we run pytrim using the call-graph-based detector to identify unused dependencies. Then, we manually verify all changes and submit PRs to the respective repositories.
pytrim removes unused dependencies in 39 projects, resulting in 39 submitted PRs. At the time of writing, 14 PRs have been merged by the project maintainers, 19 are still under review, and 6 were closed due to project-specific policies.
As an example consider the softlayer-python project (Fig. 1). pytrim correctly identified that the prettytable dependency is unused. Then, it generated a PR to remove the dependency from four configuration files (setup.py and three separate requirements.txt files). pytrim also identified the unused package name in the README.rst file and flagged it for manual review. The resulting PR, with the manual edit for the README, was merged by the project maintainers. Merged PRs to the 3 most popular projects are shown in Table II. The complete list is available in our GitHub repository.
@@
TABLE II: Merged PRs to the most popular projects (by stars).
Project PR ID Stars
napalm-automation/napalm #2243 2,374
pycqa/prospector #781 2,035
django-json-api/django-rest-framework-json-api #1285 1,238
@@ V Related Work
Existing Python tools focus primarily on detection and can be broadly categorized into two groups. The first consists of source code linters, such as autoflake [9] and pylint [10], that identify unused imports directly within .py files. While some of these, such as autoflake, can remove the identified import lines, their scope is limited to the source code; they do not modify project-level configuration files. The second group operates at the project level; tools such as deptry [12] and FawltyDeps [11] compare declared dependencies against source code imports to find unused packages, but they do not automate the removal of the corresponding import statements.